[figure: 9.7 cm]

INFLUENCE OF LOCAL ENVIRONMENT ON THE ELECTRONIC STRUCTURE AND MAGNETIC PROPERTIES

OF Ni_{1−x}Cu_{x}MnSb ALLOYS

A. Jezierski^{a} and A. Szytuła^{b}

aInstitute of Molecular Physics, Polish Academy of Sciences, 60–179 Poznań, Poland bSolid State Physics Department, Institute of Physics of the Jagiellonian University,

Reymonta 4, 30–059 Kraków, Poland

The magnetic ordering in Ni_{1−x}Cu_{x}MnSb alloys changes from ferromagnetic (x=0) to antiferromagnetic (x=1). The X–ray and neutron diffraction have shown that the systems crystallise in C1_{b} type structure. We present an influence of the local chemical ordering on the electronic and magnetic properties of Ni_{1−x}Cu_{x}MnSb alloys. The band structure and the magnetic moments were calculated by the spin–polarised LMTO–ASA method for the experimental values of the lattice parameters. The magnetic moment of Mn increases from 3.8 μ_{B} for x=0 to 4.1 μ_{B} for x=1. The magnetic moment of Ni changes from 0.26 to 0.33 μ_{B}.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : A. Jezierski

Address for correspondence : Institute of Molecular Physics Polish Academy of Sciences Smoluchowskiego 17, 60–179 Poznań

Email address : jeziersk@ifmpan.poznan.pl